OpenStudy (anonymous):

Biologists want to set up a station to test alligators in the lake for West Nile Virus. Suppose that the costs for such a station are $2,500 for setup costs and $3.00 to administer each test. You can find the average cost per animal by dividing total costs by number of animals. Write the expression that gives the average cost per animal.

OpenStudy (anonymous):

I have several parts to this question I would like help with them all. please And I would like the best answer that you have

OpenStudy (anonymous):

Total Cost = setup cost + 3 each test Total C = 2,500 + 3* t Each animal is tested once, so number of animals = number of tests = t Average cost = Total C / t Average cost = ( 2,500 + 3 * t ) / t

OpenStudy (anonymous):

c. Find the average cost per animal for 10 animals, 100 animals, and 1,000 animals.

OpenStudy (anonymous):

d. As the number of animals tested increases, what happens to the average cost to test the animals? Would the average cost ever fall below $3.00? If so, identify a value that supports your answer. If not, explain how you know.

OpenStudy (anonymous):

Here are the two that are left for this one question, then I have one more that I need help on thanks a lot for your help.

OpenStudy (anonymous):

Average cost = ( 2,500 + 3 * t ) / t Average cost (10) = ( 2,500 + 3 * 10 ) / 10 = (2,500 + 30 ) /10 = 2,530/10 = 253 Average cost (100) = ( 2,500 + 3 * 100 ) / 100 = 2,800/100 = 28 Average cost (1000) = ( 2,500 + 3 * 1000 ) / 1000 = 5,500 / 1000 = 5,5

OpenStudy (anonymous):

As the number of animals increase, we can clearly see the average cost decreases. Average cost < 3 Average cost = ( 2,500 + 3 * t ) / t < 3 ( 2,500 + 3 * t ) / t < 3 ( 2,500 + 3 * t ) < 3 * t There's no way 3 times t ( a positive number of animals) + 2,500 will ever be smaller than 3 times t. So, no.

OpenStudy (anonymous):

Thanks I hope that I have the right answers.

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

e. How many animals should be tested for the average cost to be $5.00 per animal?

OpenStudy (anonymous):

3. To estimate animal populations, biologists count the total number of animals in a small section of a habitat. The total population of animals is directly proportional to the size of the habitat (in acres) polled.